The length of a rectangle exceeds the width by 7. If the area is 60cm^2 find the length of the rectangle

Answer by jorel555(1290)

Let w be the width of the rectangle. Then the length is w+7. So:
w(w+7)=60
w²+7w-60=0
(w+12)(w-5)=0
w=-12 or 5
Throwing out tne negative result, we get the width to be 5 cm, and the length to be 12 cm. ☺☺☺☺